OpenStudy (anonymous):

Find the first five terms of the sequence a_n=1/3^n

OpenStudy (anonymous):

a1=1/3^1=1/3, a2=1/3^2=1/9 a3=1/3^3=1/27 a4=1/3^4=1/81 a5=1/3^5=1/243 You could try to do this on your own through this: \[\sum_{n=1}^{5}(1/3)^{n}\]

OpenStudy (anonymous):

THANKS